# Poncelet triangles: two harmonious loci and two attractive envelopes

### v1

@@ 1. Introduction
Consider a Poncelet porism of triangles [3] inscribed in a first ellipse $\mathcal{E}$ and circumscribing a second, nested one called $\mathcal{E}_{c}$, Figure˜1 (left). We have been interested in loci of triangle centers over Poncelet since we first encountered [21, 13], and are motivated by the continued exploration of related kinematic phenomena, e.g., [4, 9, 11].
@@
  • Section˜4: in [17] it is shown that if $\mathcal{E}$ is a circle, the locus of the isogonal conjugate – first studied by J. Neuberg and E. Lemoine in the 1880s, Figure˜4 (left) – of a fixed point with respect to Poncelet triangles is also a circle, Figure˜1 (right). Here we let $\mathcal{E}$ be a generic ellipse, and find that the same locus – note that isogonal conjugation is not affinely invariant – is a conic, i.e., a degenerate degree-four curve. Interestingly, this conic becomes a line if $P$ is on $\mathcal{E}$, and to a parabola if it lies on the degree-4 boundary of the region swept by the circumcircle, Section˜4.1.
@@
Our proofs are based on analytic geometry, often producing long, explicit expressions for the loci or envelopes with the aid of a Computer Algebra System (CAS). The reader is encouraged to replace them for a more concise argument based on the tools of algebraic geometry descrbied in [21] and deployed in [15, 16].
@@ 2. Symmetric parametrization
Identifying $\mathbb{R}^{2}$ with $\mathbb{C}$, consider the following parameterization for Poncelet triangles inscribed in $\mathbb{T}$, the unit circle centered at the origin, as derived in [6, Def. 3] and based on the work in [2] on Blaschke products:
@@
This is generalized to a Poncelet triangle family $\mathcal{T}$ interscribed between any two nested ellipses $\mathcal{E},\mathcal{E}_{c}$ by applying an affine transformation that sends $\mathbb{T}$ to $\mathcal{E}$. Let $z_{1},z_{2},z_{3}\in\mathcal{E}$ be the varying vertices of the Poncelet triangles. The statements below are reproduced from [5, Sec.2.2], where their proofs can also be found:
@@ 4. Isogonal locus
In particular, isogonal conjugation sends a triangle’s circumcircle to the line at infinity (and vice-versa) [20, Isogonal Conjugate]. If the barycentrics of $P$ are $[z_{1}:z_{2}:z_{3}]$, $P^{\dagger}=[l_{1}^{2}/z_{1}:l_{2}^{2}/z_{2}:l_{3}^{2}/z_{3}]$, where the $l_{i}$ are the sidelengths. An alternative construction is shown in Figure˜4, based on the shared circumcircle of the pedal triangles of $P,P^{\dagger}$.
@@
An expression for the complex circumcenter $O$ of a triangle with vertices $\{x,y,z\}$ in terms of a ratio of determinants is given in [1, Lemma 6.24, p.108]:
@@ 4.2. Conic type
Figure 8. $P$ interior to $\mathcal{E}_{c}$, elliptic locus of $P^{\dagger}$ is doesn’t intersect $\mathcal{E}$. The green region is $\mathcal{R}$, i.e., what is swept by the circumcircle of $ABC$ over Poncelet. left: $P$ at a generic position within $\mathcal{E}_{c}$ right: when that point is a focus of $\mathcal{E}_{c}$, $P^{\dagger}$ is stationary on the other focus since the foci of an inconic are a conjugate pair [12].
@@ 5. Two envelopes with a circular caustic
Let $\mathcal{T}_{\circ}$ denote the Poncelet family inscribed in $\mathcal{E}$ (semiaxes $a,b$) and circumscribing a circular caustic $\mathcal{K}=(C,r)$ with $C=[x_{c},y_{c}]$. The pair $\mathcal{E},\mathcal{K}$ will admit a family of Poncelet triangles if $r$ is given by [5, Prop.2]:
@@
with $c^{2}=a^{2}-b^{2}$. Let $\mathcal{L}_{3}$ denote the locus of the circumcenter $X_{3}$ over $\mathcal{T}_{\circ}$, which must be a conic since it is a fixed (and trivial) linear combination of $X_{2}$ and $X_{3}$ [6, Thm.2].
In [5, Prop.7] we prove that the foci $F_{3}$ and $F_{3}^{\prime}$ of $\mathcal{L}_{3}$ are collinear with $C$, with each lying on an axis of $\mathcal{E}$ at:
@@
Expressions for the semi-axis lengths $a_{3},b_{3}$ of $\mathcal{L}_{3}$ are also derived in [5, Prop.7]:
@@ 5.1. Circumcircle
Let $C=[x_{c},y_{c}]$ be the center of caustic and $r$ given by Equation˜6; the incircle $\mathcal{K}$ is then given by $(x-x_{c})^{2}+(y-y_{c})^{2}=r^{2}$. Let $\lambda=\cos{u}+i\sin{u}$ be of the symmetric parametrization of Poncelet triangles as in [7, Lem.1]. Obtain the circumcircle $\mathcal{K}^{\prime}$ as a function of $u$:
@@
$\mathcal{K}_{1}$ touches $\mathcal{E}$ at [±$\left[\pm(a/b)\sqrt{b^{2}-y_{c}^{2}},y_{c}\right]$; $\mathcal{K}_{2}$ touches $\mathcal{E}$ at $\left[x_{c},\pm(b/a)\sqrt{a^{2}-x_{c}^{2}}\right]$. ∎
@@
Given two fixed nested (resp. unnested) circles, the locus of the center of circles simultaneously tangent to them is a conic with foci on their centers and with major axis length equal to the difference (resp. sum) of their radii [14, thm.14, p.57]. In our case, the circumcircles of a Poncelet family circumscribing a circle are tangent to the two circular boundaries of their envelope. It follows from Equation˜8:
@@ 5.2. Incircle-circumcircle radical axis
Let $\Gamma$ denote the radical axis of incircle and circumcircle of a triangle, i.e., the line whose points have the same power with respect to both circles [20, Radical line]. $X_{1}X_{3}$ is perpendicular to ${\Gamma}$ since it is the line that passes through both circle centers. Incidentally, the intersection of $X_{1}X_{3}$ with ${\Gamma}$ is named $X_{3660}$ on [10].
@@
By continuity, if $C$ is on $\mathcal{E}_{\triangle}$, $\mathcal{T}_{\circ}$ will contain an equilateral, and is henceforth called $\mathcal{T}_{\triangle}$. In [5, Prop.9], it is shown that while over $\mathcal{T}_{\circ}$, the locus of $X_{36}$ (the circumcircle-inverse of the incenter $X_{1}$) sweeps a circle, over $\mathcal{T}_{\triangle}$ it degenerates to the line:
@@
The Feuerbach point $X_{11}$ is where incircle and nine-point circle touch. Therefore, the radical axis of said circles is tangent to the incircle at $X_{11}$. Let $L_{101}$ denote said axis as in the table at the end of [20, Radical line]. Therefore:
@@
In [5, Prop.26] we show that over $\mathcal{T}_{\triangle}$, $X_{11}$ is stationary. Therefore, in such a case said envelope is undefined.
[…]
@@ Appendix B Triangle centers
$X_{2}$ barycenter/centroid 1::1::1 : : meet of medians
@@
Table 2. Kimberling codes for various triangle centers mentioned here, along with their names, barycentric coordinates (only the first shown, the other two can be obtaine by cyclical replacement), and construction notes [10]. The isogonal conjugate of a point with barycentrics $[z_{1}:z_{2}:z_{3}]$ is $[l_{1}^{2}/z_{1}:l_{2}^{2}/z_{2}:l_{3}^{2}/z_{3}]$, where the $l_{i}$ are the sidelengths.
@@ Appendix D Circumcircle envelope
The envelope of the region $\partial\mathcal{R}$ swept by the circumcircle is given by the zeros of the following expression (of degree four), where $f,g$ are the foci of the Poncelet caustic of the $\mathbb{T}$-inscribed family, and $a,b$ are the semi-axis’ lengths of $\mathcal{E}$:
$\partial\mathcal{R}(a,b,f,g,x,y)=-b^{6}\left[(-1+f_{y}^{2})\,g_{x}^{2}+(-1+f_{y}g_{y})^{2}+f_{x}^{2}(-1+g_{x}^{2}+g_{y}^{2})\right]x^{2}$

### v3

# Poncelet triangles: conic loci of the orthocenter and of the isogonal conjugate of a fixed point

@@ 1. Introduction
Consider a Poncelet porism of triangles [5] inscribed in a first ellipse $\mathcal{E}$ and circumscribing a second, nested one called $\mathcal{E}_{c}$, Figure˜1 (left). We have been interested in loci of triangle centers over Poncelet since we first encountered [25, 16], and are motivated by the continued exploration of related kinematic phenomena, e.g., [6, 11, 13, 14].
@@
  • Section˜4: referring to Figure˜1 (right), in [21] it is shown that if $\mathcal{E}$ is a circle, the isogonal conjugate [^1] A type of inversive transformation first studied by J. Neuberg and E. Lemoine in the 1880s., – see Figure˜4 (left) and [1, 4, 20]– of a fixed point with respect to Poncelet triangles sweeps a circle, Figure˜1 (right). Here we let $\mathcal{E}$ be a generic ellipse, and find that the same locus – note that isogonal conjugation is not affinely invariant – is a conic. Interestingly, said locus degeneratesto a parabola if it lies on the degree-4 boundary of the region swept by the circumcircle, Section˜4.1, and to a line if $P$ lies on $\mathcal{E}$. Furthermore, over all $P$ on $\mathcal{E}$, said line-loci envelop an ellipse, concentric with $\mathcal{E}_{c}$.
@@
Our proofs are based on analytic geometry, often producing long, explicit expressions for the loci or envelopes with the aid of a Computer Algebra System (CAS). The reader is encouraged to replace them for a more concise argument based on the tools of algebraic geometry described in [25] and deployed in [18, 19].
@@ 2. Symmetric parametrization
Identifying $\mathbb{R}^{2}$ with $\mathbb{C}$, consider the following parameterization for Poncelet triangles inscribed in $\mathbb{T}$, the unit circle centered at the origin, as derived in [8, Def. 3] and based on the work in [3] on Blaschke products:
@@
This is generalized to a Poncelet triangle family $\mathcal{T}$ interscribed between any two nested ellipses $\mathcal{E},\mathcal{E}_{c}$ by applying an affine transformation that sends $\mathbb{T}$ to $\mathcal{E}$. Let $z_{1},z_{2},z_{3}\in\mathcal{E}$ be the varying vertices of the Poncelet triangles. The statements below are reproduced from [7, Sec.2.2], where their proofs can also be found:
@@ 4. Isogonal locus
In particular, isogonal conjugation sends a triangle’s circumcircle to the line at infinity (and vice-versa) [24, Isogonal Conjugate]. If the barycentrics of $P$ are $z_{1}:z_{2}:z_{3}]$, $P^{\dagger}=[l_{1}^{2}/z_{1}:l_{2}^{2}/z_{2}:l_{3}^{2}/z_{3}]$, where the $l_{i}$ are the sidelengths. An alternative construction is shown in Figure˜4, based on the shared circumcircle of the pedal triangles of $P,P^{\dagger}$. For more information, see [4, 20].
@@
An expression for the complex circumcenter $O$ of a triangle with vertices $\{x,y,z\}$ in terms of a ratio of determinants is given in [2, Lemma 6.24, p.108]:
@@ 4.2. Conic type
Figure 8. $P$ interior to $\mathcal{E}_{c}$, elliptic locus of $P^{\dagger}$ is doesn’t intersect $\mathcal{E}$. The green region is $\mathcal{R}$, i.e., what is swept by the circumcircle of ABCABC over Poncelet. left: $P$ at a generic position within $\mathcal{E}_{c}$ right: when that point is a focus of $\mathcal{E}_{c}$, $P^{\dagger}$ is stationary on the other focus since the foci of an inconic are a conjugate pair [15].
@@ 5. Two envelopes with a circular caustic
Let $\mathcal{T}_{\circ}$ denote the Poncelet family inscribed in $\mathcal{E}$ (semiaxes $a,b$) and circumscribing a circular caustic $\mathcal{K}=(C,r)$ with $C=[x_{c},y_{c}]$. The pair $\mathcal{E},\mathcal{K}$ will admit a family of Poncelet triangles if $r$ is given by [7, Prop.2]:
@@
with $c^{2}=a^{2}-b^{2}$. Let $\mathcal{L}_{3}$ denote the locus of the circumcenter $X_{3}$ over $\mathcal{T}_{\circ}$, which must be a conic since it is a fixed (and trivial) linear combination of $X_{2}$ and $X_{3}$ [8, Thm.2].
In [7, Prop.7] we prove that the foci $F_{3}$ and $F_{3}^{\prime}$ of $\mathcal{L}_{3}$ are collinear with $C$, with each lying on an axis of $\mathcal{E}$ at:
@@
Expressions for the semi-axis lengths $a_{3},b_{3}$ of $\mathcal{L}_{3}$ are also derived in [7, Prop.7]:
@@ 5.1. Circumcircle
Let $C=[x_{c},y_{c}]$ be the center of caustic and $r$ given by Equation˜6; the incircle $\mathcal{K}$ is then given by $(x-x_{c})^{2}+(y-y_{c})^{2}=r^{2}$. Let $\lambda=\cos{u}+i\sin{u}$ be of the symmetric parametrization of Poncelet triangles as in [9, Lem.1]. Obtain the circumcircle $\mathcal{K}^{\prime}$ as a function of $u$:
@@
$\mathcal{K}_{1}$ touches $\mathcal{E}$ at $\left[\pm(a/b)\sqrt{b^{2}-y_{c}^{2}},y_{c}\right]$; $\mathcal{K}_{2}$ touches $\mathcal{E}$ at $\left[x_{c},\pm(b/a)\sqrt{a^{2}-x_{c}^{2}}\right]$. ∎
@@
Given two fixed nested (resp. unnested) circles, the locus of the center of circles simultaneously tangent to them is a conic with foci on their centers and with major axis length equal to the difference (resp. sum) of their radii [17, thm.14, p.57]. In our case, the circumcircles of a Poncelet family circumscribing a circle are tangent to the two circular boundaries of their envelope. It follows from Equation˜8:
@@ 5.2. Incircle-circumcircle radical axis
Let $\Gamma$ denote the radical axis of incircle and circumcircle of a triangle, i.e., the line whose points have the same power with respect to both circles [24, Radical line]. $X_{1}X_{3}$ is perpendicular to $\Gamma$} since it is the line that passes through both circle centers. Incidentally, the intersection of $X_{1}X_{3}$ with $\Gamma$} is named $X_{3660}$ on [12].
@@
By continuity, if $C$ is on $\mathcal{E}_{\triangle}$, $\mathcal{T}_{\circ}$ will contain an equilateral, and is henceforth called $\mathcal{T}_{\triangle}$. In [7, Prop.9], it is shown that while over $\mathcal{T}_{\circ}$, the locus of $X_{36}$ (the circumcircle-inverse of the incenter $X_{1}$) sweeps a circle, over $\mathcal{T}_{\triangle}$ it degenerates to the line:
@@
The Feuerbach point $X_{11}$ is where incircle and nine-point circle touch. Therefore, the radical axis of said circles is tangent to the incircle at $X_{11}$. Let $L_{101}$ denote said axis as in the table at the end of [24, Radical line]. Therefore:
@@
In [7, Prop.26] we show that over $\mathcal{T}_{\triangle}$, $X_{11}$ is stationary. Therefore, in such a case said envelope is undefined.
The orthic axis, called $L_{3}$ in (resp. anti-orthic axis, $L_{1}$) in [24, Radical line], is the radical axis of the circumcircle and the nine-point (resp. Bevan) circle. The latter is centered on $X_{5}$ (resp. $X_{40}$) and has radius $R/2$ (resp. $2R$).
@@ Appendix B Triangle centers
$X_{2}$ barycenter/centroid $1::$ meet of medians
@@
Table 2. Kimberling codes for various triangle centers mentioned here, along with their names, barycentric coordinates (only the first shown, the other two can be obtaine by cyclical replacement), and construction notes [12]. The isogonal conjugate of a point with barycentrics $z_{1}:z_{2}:z_{3}]$ is $l_{1}^{2}/z_{1}:l_{2}^{2}/z_{2}:l_{3}^{2}/z_{3}]$, where the $l_{i}$ are the sidelengths.
@@ Appendix D Circumcircle envelope
The envelope of the region $\partial\mathcal{R}$ swept by the circumcircle is given by the zeros over $x,y$ of:
$-b^{6}\left[(-1+f_{y}^{2})\,g_{x}^{2}+(-1+f_{y}g_{y})^{2}+f_{x}^{2}(-1+g_{x}^{2}+g_{y}^{2})\right]x^{2}$